1. Solve the following simultaneous equations for x and y.

1.1. 7x+2y=1 and 4x+3y=8
1.2. x+y=5 and x-y=3

1.3 x+3y=-30 and y=3x
1.4. 2x+5y=9 and x+3y=5

1.5. (a+b)/3=2+2b and (2a-4b)/5= 23/5-b
1.6. y=2x+3 and y=-3x–7

1.7.x+2y+3 =7x–y and 2x–3y+16 =7x–y
1.8. (x+y)/(3-y)=1/2 and (x-y)/(3+2y)=4

Answer by Alan3354(69443)

1. Solve the following simultaneous equations for x and y.

1.1. 7x+2y=1 and 4x+3y=8
Multiply the 1st eqn by 3, the 2nd by 2, then subtract.
------------
x+y=5
x-y=3
--------------- Add
=======================
x+3y=-30 and y=3x
Sub for y in the 1st eqn.
-------------
2x+5y=9
x+3y=5 --> x = -3y+5
Sub for x in the 1st eqn
---
Or, multiply the 2nd eqn by 2, then subtract.
==========================================

(a+b)/3=2+2b and (2a-4b)/5= 23/5-b
Eliminate fractions.
(a+b)/3=2+2b times 3 --> a+b = 6 + 6b
(2a-4b)/5= 23/5-b times 5 --> 2a-4b = 23 - 5b
Use one of the methods above.
===================================
y=2x+3 and y=-3x–7
Sub for y
========================
x+2y+3 =7x–y and 2x–3y+16 =7x–y
Collect terms, use a method from above.
=============================
(x+y)/(3-y)=1/2 and (x-y)/(3+2y)=4
Eliminate fractions, collect terms, etc.